# Infinite-component $BF$ topological field theory: Nexus of fracton order, Toeplitz braiding, and non-Hermitian Amplification

### v2

@@ Abstract
Building on the infinite-component Chern–Simons theory of three-dimensional fracton phases by Ma et al. [Phys. Rev. B 105, 195124 (2022)] and the Toeplitz braiding of anyons by Li et al. [Phys. Rev. B 110, 205108 (2024)], we show that stacking (3+1)(3+1)D BFBF topological field theories along a fourth spatial direction gives rise to an exotic class of four-dimensional fracton phases. Their low-energy physics is governed by a new field-theoretic framework—infinite-component BFBF (iBFBF) theories—characterized by asymmetric integer Toeplitz $K$ matrices. Under open boundary conditions, iBFBF theories exhibit a striking phenomenon: Toeplitz particle–loop braiding, where a particle and a loop placed on opposite three-dimensional boundaries acquire a finite braiding phase even at infinite separation. This nonlocal braiding admits a geometric interpretation: transporting the particle induces a winding boundary trajectory on the opposite boundary that encircles the loop. We show that this robustness originates from boundary zero singular modes (ZSMs) of Toeplitz $K$ matrices revealed by singular value decomposition (SVD), rather than from the eigenvalue zero modes responsible for previously known Toeplitz braiding of anyons. We analytically and numerically study representative iBFBF theories with Hatano–Nelson–type and non-Hermitian Su–Schrieffer–Heeger–type $K$ matrices, establishing a universal correspondence between ZSMs and Toeplitz particle–loop braiding. Our results identify boundary zero singular modes as the mechanism behind Toeplitz particle–loop braiding and establish infinite-component $BF$ theory as a predictive framework for higher-dimensional fracton topological orders.
@@ I.1 Infinite-component Chern-Simons theory, anyonic Toeplitz braiding, and topological edge states
Recently, fracton topological order (FTO) has emerged as a distinctive frontier in quantum many-body systems [1, 2, 3, 4, 5], driving advances in condensed matter and quantum physics [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Solvable realizations of FTO, known as fracton codes, include the 3d [^1] In this paper, “(n+1)(n+1)D” refers to $(n+1)$-dimensional spacetime with $n$-dimensional real space. We avoid using “nnD” or “(n)(n)D” unless otherwise specified. When referring specifically to spatial dimensions, we use “nnd”, e.g., 33d ground state, 33d topological order, 33d cubic lattice and 11d closed string. X-cube model and Haah’s code [3, 4, 5]. In arbitrary dimensions, generalizations called “tetradigit stabilizer codes” have been proposed [13, 14, 15, 16]. Their excitations—collectively referred to as fractonic excitations—exhibit restricted mobility: point-like fractons are immobile, lineons move only along lines, and planeons are confined to planes. This sharply contrasts with liquid-like topological orders such as the toric code, where anyons propagate freely via string operators. Studies of FTO have further inspired subsystem symmetry-protected topological phases [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] and many-body systems with conserved multipole moments [34, 35, 36]. The former generalize global symmetries to lower-dimensional subsystems; the latter impose strong kinematic constraints, yielding unconventional quantum phenomena [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95].
@@
A large subclass of FTOs exhibits a foliation structure, allowing them to be viewed as stacks of coupled 2d topological orders. This perspective provides a natural route to field-theoretical descriptions of fracton phases. Recent studies have developed stacking-based continuum or effective field theories [96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110]. Among these, the infinite-component Chern–Simons (iCS) theory—constructed by stacking multi-component Chern–Simons layers—has proven particularly promising [103, 104, 105, 106, 107, 108, 109, 110]. Before proceeding, we briefly recall the basic structure of the (2+1)(2+1)D multi-component Chern–Simons theory, which serves as a prototypical topological field theory capturing the low-energy physics of Abelian topological orders.
@@
In (2+1)(2+1)D multi-component Chern–Simons theory, topological quantities such as braiding statistics, ground-state degeneracy on a torus, and chiral central charge are encoded in the symmetric integer $K$ matrix [111, 112]. This framework covers many topologically ordered phases, including multilayer fractional quantum Hall states and certain spin liquids [112]. It also provides a systematic method for classifying symmetry-enriched (SET) and symmetry-protected (SPT) phases in 2d [113, 114, 115, 116, 117, 118, 119]. The Abelian Chern–Simons theory considered here follows a hydrodynamical approach, distinct from the flux-attachment formulation [120, 121, 122, 123].
@@
To construct 3d FTOs, we treat each 2d ($xy$ plane) Abelian topological order described by (2+1)(2+1)D multi-component Chern–Simons theory as a 2d “layer” and stack them along the $z$ direction, coupling adjacent layers via additional Chern–Simons terms that preserve translational invariance. This yields a multi-component Chern–Simons theory whose topological data are encoded in a symmetric integer block-Toeplitz ($K$) matrix [124, 110]. The action of such a theory is
@@
Here $a^{I,J}$ ($I,J=1,\ldots,N$) are $U(1)$ compact gauge fields, and implicit summation is assumed over these indices, which also label positions along $z$. The absence of lower-left and upper-right blocks in $K$ indicates open boundary conditions (OBC) along the stacking direction. In the limit $N\to\infty$, this construction yields the infinite-component Chern–Simons (iCS) theory. The system boundaries fall into two classes: the $xy$ surfaces of each layer and the $z$ boundaries along the stacking direction. While the $xy$ boundaries are conventional and can be described by stacked conformal field theories, the novel $z$ boundaries were only understood after the discovery of Toeplitz braiding in Ref. [110].
@@
Figure 1: Illustration of stacking and Toeplitz particle–loop braiding in iBF theories. Each so-called “layer” is a genuine three-dimensional BFBF theory defined on a three-torus and is schematically represented by a light-blue cube. These 3d layers are stacked along the $w$ direction (i.e., the extra spatial dimension) and coupled by interlayer $BF$ terms (wavy lines), giving rise to two (3+1)(3+1)D boundaries labeled as $w_{1}$ and $w_{2}$. Panel (a) shows Toeplitz particle–loop braiding between a loop excitation (red) on the $w_{1}$ boundary and a particle excitation (red) on the $w_{2}$ boundary. A nontrivial braiding phase arises when the particle (red) is adiabatically transported such that its induced boundary trajectory (blue) on the opposite boundary—namely, the projected “shadow” of the particle (red) onto the $w_{1}$ boundary—winds around the loop (red). This process is encoded in the lower-left elements of $K^{-1}$, highlighted in panel (c). Panel (b) depicts the opposite configuration, encoded in the upper-right elements of $K^{-1}$ highlighted in panel (d). Panel (e) summarizes the correspondence between the boundary localization of left and right zero singular modes (LZSMs and RZSMs), the structure of $K^{-1}$, and the resulting Toeplitz particle–loop braiding. These summarized results are higher-dimensional generalization of Toeplitz braiding introduced in Ref. [110].
@@
The usual (3+1)(3+1)D multi-component $BF$ theories [126, 127] provide a versatile framework for describing 3d topological orders. Analogous to the (2+1)(2+1)D Chern–Simons description of 2d systems, when pointlike and loop excitations are represented as charges and fluxes of an Abelian finite gauge group, the $BF$ theory and its twisted variants [123, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137] form a unified field-theoretic scheme for capturing 3d topological orders and gauged SPT phases. A finite set of canonical $BF$ terms $BdA$ (with 2-form $B$ and 1-form $A$ fields), together with twisted terms such as $AAdA$, AAAAAAAA, AABAAB, BBBB, and the $\theta$-term $dAdA$, encodes rich topological data—including particle–loop, multi-loop, and Borromean rings braiding [127, 138, 139, 140, 141, 142, 143, 130, 144, 145, 146, 147, 148, 149, 150, 151, 134], emergent fermionic statistics [152, 136, 131, 153], and topological responses [154, 155, 129, 123, 156, 137, 157, 158]. Refs. [159, 160, 161] further showed how global symmetries fractionalize on loop excitations, introducing “mixed three-loop braiding”, which underlies the classification of higher-dimensional SET phases. These braiding processes obey stringent consistency conditions [135] ensuring gauge invariance and allowing for non-Abelian fusion and shrinking rules even when loop charges remain Abelian [134, 162]. Building on this foundation, a diagrammatic representation for higher-dimensional topological orders—including 4d cases [163]—was recently developed in Ref. [164].
To this end, we generalize the above established multi-component $BF$ framework to an iBFBF theory—to describe 4d FTOs and Toeplitz braiding phenomena. In analogy to the iCS construction, we couple multi-component (3+1)(3+1)D BFBF theories to formulate an effective field theory for 4d FTOs: $S=\int\frac{A_{I,J}}{2\pi}\,b^{I}\wedge da^{J}$. Stacking $N$ layers of such (3+1)(3+1)D field theories and coupling neighboring 3d “layers” via interlayer $BF$ terms while preserving translational symmetry yields a multi-component $BF$ theory characterized by a Toeplitz $K$ matrix with integer elements:
S=KI,J2πbIdaJ,
=K(ACBACBACBACBA)
@@
Given an iBFBF theory with nearest-neighbor coupling, a natural question arises: under what conditions does nonlocal braiding along the stacking direction emerge? Is there an analogue of boundary-mode-induced nonlocal braiding—i.e., Toeplitz braiding— along the stacking direction? To address this, we classify iBFBF theories into two types. The first involves symmetric $K$ matrices, where $C=B^{\mathsf{T}}$ in Eq. (2); in this case, the Toeplitz braiding analysis parallels that of iCS theories [110], except that the particle–particle braiding of Ref. [110] now corresponds to particle–loop braiding in the $BF$ context.
@@
When all “left” ZSMs (LZSMs) localize on one $w$ boundary and “right” ZSMs (RZSMs) on the other [^4] Here the “left” and the “right” label internal sectors of the SVD rather than spatial sides., either the lower-left or upper-right elements of $K^{-1}$ decay exponentially with system size. Consequently, nontrivial braiding arises when particles and loops occupy specific opposite $w$ boundaries, as illustrated in Figs. 1(a)–1(b). The corresponding braiding phases are encoded in the lower-left or upper-right elements of $K^{-1}$ [Figs. 1(c)–1(d)]. In either case, exchanging the particle and loop positions rapidly suppresses the braiding phase upon increasing the size of the $w$ direction. In particular, iBFBF theories with $K$ matrices resembling the nSSH Hamiltonians can exhibit nontrivial Toeplitz braiding in both configurations shown schematically in Figs. 1(a) and 1(b), as the $K$ matrices may host two LZSMs (RZSMs) that localize at distinct $w$ boundaries. The relation between the locations of the LZSMs and RZSMs, the patterns of $K^{-1}$, and the resulting braiding statistics is summarized in Fig. 1(e).
@@
The paper is organized as follows. Sec. II establishes the braiding structure of multi-component $BF$ theories and introduces the stacking construction that gives rise to iBFBF field theories, emphasizing why non-diagonal Toeplitz $K$ matrices are essential. Sec. III and Sec. IV uncover Toeplitz braiding in iBFBF theories whose $K$ matrices inherit the mathematical structures of the HN and nSSH Hamiltonians, respectively. These sections show how SVD sharply exposes boundary zero singular modes and reveal a direct analogy to non-Hermitian directional amplification. Sec. V closes with a brief summary and outlook.

### v3

# Infinite-component $BF$ topological field theory: connection of fracton order, Toeplitz braiding, and non-Hermitian Amplification

@@ Abstract
Building on the infinite-component Chern–Simons theory of three-dimensional fracton phases by Ma et al. [Phys. Rev. B 105, 195124 (2022)] and the Toeplitz braiding of anyons by Li et al. [Phys. Rev. B 110, 205108 (2024)], we show that stacking (3+1)(3+1)D BFBF topological field theories, which serve as low-energy effective descriptions of a class of three-dimensional topological orders, along a fourth spatial direction gives rise to an exotic class of four-dimensional fracton phases. Their low-energy physics is governed by a new field-theoretic framework, namely infinite-component BFBF (iBFBF) theories, characterized by asymmetric integer Toeplitz $K$ matrices. Under open boundary conditions along the stacking direction, iBFBF theories with properly chosen $K$ matrices exhibit a striking phenomenon termed Toeplitz particle–loop braiding, where a particle and a loop placed on opposite three-dimensional boundaries acquire a strongly oscillating yet robustly nonvanishing braiding phase even at infinite separation. This nonlocal braiding admits a geometric interpretation: adiabatically transporting the particle induces a winding boundary trajectory on the opposite boundary that encircles the loop. We show that this robustness originates from boundary zero singular modes (ZSMs) of Toeplitz $K$ matrices revealed by singular value decomposition, rather than from boundary zero eigenmodes responsible for previously known Toeplitz braiding of anyons, and that the same ZSM mechanism also underlies directional amplification in the rapidly developing field of non-Hermitian physics. We analytically and numerically study representative iBFBF theories with Hatano–Nelson–type and non-Hermitian Su–Schrieffer–Heeger–type $K$ matrices, establishing a universal correspondence between ZSMs and Toeplitz particle–loop braiding. Our results identify boundary zero singular modes as the operative mechanism behind Toeplitz particle–loop braiding and establish infinite-component $BF$ theory as a predictive framework for higher-dimensional fracton topological orders.
@@ I.1 Infinite-component Chern-Simons theory, anyonic Toeplitz braiding, and topological edge states
Recently, fracton topological order (FTO) has emerged as a distinctive frontier in quantum many-body systems [1, 2, 3, 4, 5], driving advances in condensed matter and quantum physics [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Solvable realizations of FTO, known as fracton codes, include the 3D [^1] In this paper, “(n+1)(n+1)D” refers to $(n+1)$-dimensional spacetime with $n$-dimensional real space. When referring specifically to spatial dimensions, we use “nnD”. X-cube model and Haah’s code [3, 4, 5]. In arbitrary dimensions, generalizations called “tetradigit stabilizer codes” have been proposed [13, 14, 15, 16]. Their excitations—collectively referred to as fractonic excitations—exhibit restricted mobility: point-like fractons are immobile, lineons move only along lines, and planeons are confined to planes. This sharply contrasts with liquid-like topological orders such as the toric code, where anyons propagate freely via string operators. Studies of FTO have further inspired subsystem symmetry-protected topological phases [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] and many-body systems with conserved multipole moments [34, 35, 36]. The former generalize global symmetries to lower-dimensional subsystems; the latter impose strong kinematic constraints, yielding unconventional quantum phenomena [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95].
@@
A large subclass of FTOs exhibits a foliation structure, allowing them to be viewed as stacks of coupled 2D topological orders. This perspective provides a natural route to field-theoretical descriptions of fracton phases. Recent studies have developed stacking-based continuum or effective field theories [96, 97, 98, 99, 100, 101, 17, 102, 103, 104, 105, 106, 107, 108, 109]. Among these, the infinite-component Chern–Simons (iCS) theory—constructed by stacking multicomponent Chern–Simons layers—has proven particularly promising [102, 103, 104, 105, 106, 107, 108, 109]. Before proceeding, we briefly recall the basic structure of the (2+1)(2+1)D multicomponent Chern–Simons theory, which serves as a prototypical topological field theory capturing the low-energy physics of Abelian topological orders.
@@
In (2+1)(2+1)D multicomponent Chern–Simons theory, topological quantities such as braiding statistics, ground-state degeneracy on a torus, and chiral central charge are encoded in the symmetric integer $K$ matrix [110, 111]. This framework covers many topologically ordered phases, including multilayer fractional quantum Hall states and certain spin liquids [111]. It also provides a systematic method for classifying symmetry-enriched (SET) and symmetry-protected (SPT) phases in 2D [112, 113, 114, 115, 116, 117, 118]. The Abelian Chern–Simons theory considered here follows a hydrodynamical approach, distinct from the flux-attachment formulation [119, 120, 121, 122].
@@
To construct 3D FTOs, we treat each 2D ($xy$ plane) Abelian topological order described by (2+1)(2+1)D multicomponent Chern–Simons theory as a 2D “layer” and stack them along the $z$ direction, coupling adjacent layers via additional Chern–Simons terms that preserve translational invariance. This yields a multicomponent Chern–Simons theory whose topological data are encoded in a symmetric integer block-Toeplitz ($K$) matrix [123, 109]. The action of such a theory is
@@
Here $a^{I,J}$ ($I,J=1,\ldots,N$) are $U(1)$ compact gauge fields, and implicit summation is assumed over these indices, which also label positions along $z$. The absence of lower-left and upper-right blocks in $K$ indicates open boundary conditions (OBC) along the stacking direction. In the limit $N\to\infty$, this construction yields the iCS theory. The system boundaries fall into two classes: the $xy$ surfaces of each layer and the $z$ boundaries along the stacking direction. While the $xy$ boundaries are conventional and can be described by stacked conformal field theories, the novel $z$ boundaries were only understood after the discovery of Toeplitz braiding in Ref. [109].
@@
Figure 1: Illustration of stacking and Toeplitz particle–loop braiding in iBFBF theories. Each so-called “layer” is a genuine three-dimensional BFBF theory defined on a three-torus and is schematically represented by a light-blue cube. These 3D layers are stacked along the $w$ direction (i.e., the extra spatial dimension) and coupled by interlayer $BF$ terms (wavy lines), giving rise to two (3+1)(3+1)D boundaries labeled as $w_{1}$ and $w_{2}$. Panel (a) shows Toeplitz particle–loop braiding between a loop excitation (red) on the $w_{1}$ boundary and a particle excitation (red) on the $w_{2}$ boundary. A nontrivial braiding phase arises when the particle (red) is adiabatically transported such that its induced boundary trajectory (blue) on the opposite boundary—namely, the projected “shadow” of the particle (red) onto the $w_{1}$ boundary—winds around the loop (red). This process is encoded in the lower-left elements of $K^{-1}$, highlighted in panel (c). Panel (b) depicts the opposite configuration, encoded in the upper-right elements of $K^{-1}$ highlighted in panel (d). Panel (e) summarizes the correspondence between the boundary localization of left and right zero singular modes (LZSMs and RZSMs), the structure of $K^{-1}$, and the resulting Toeplitz particle–loop braiding. These summarized results are higher-dimensional generalization of Toeplitz braiding introduced in Ref. [109].
@@
The usual (3+1)(3+1)D multicomponent $BF$ theories [125, 126] provide a versatile framework for describing 3D topological orders. Analogous to the (2+1)(2+1)D Chern–Simons description of 2D systems, when pointlike and loop excitations are represented as charges and fluxes of an Abelian finite gauge group, the $BF$ theory and its twisted variants [122, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136] form a unified field-theoretic scheme for capturing 3D topological orders and gauged SPT phases. A finite set of canonical $BF$ terms $BdA$ (with 2-form $B$ and 1-form $A$ fields), together with twisted terms such as $AAdA$, AAAAAAAA, AABAAB, BBBB, and the $\theta$-term $dAdA$, encodes rich topological data—including particle–loop, multiloop, and Borromean rings braiding [126, 137, 138, 139, 140, 141, 142, 129, 143, 144, 145, 146, 147, 148, 149, 150, 133], emergent fermionic statistics [151, 135, 130, 152], and topological responses [124, 153, 128, 122, 154, 136, 155, 156]. Refs. [157, 158, 159] further showed how global symmetries fractionalize on loop excitations, introducing “mixed three-loop braiding”, which underlies the classification of higher-dimensional SET phases. These braiding processes obey stringent consistency conditions [134] ensuring gauge invariance and allowing for non-Abelian fusion and shrinking rules even when loop charges remain Abelian [133, 160]. Building on this foundation, a diagrammatic representation for higher-dimensional topological orders—including 4d cases [161]—was recently developed in Ref. [162]. Moreover, recent studies have explored novel self-statistics of fully immobile excitations in FTOs [163], demonstrating that braiding statistics in FTOs can differ from those in liquid-like topological orders with fully mobile excitations.
To this end, we generalize the above established multicomponent $BF$ framework to an iBFBF theory—to describe 4d FTOs and Toeplitz braiding phenomena. In analogy to the iCS construction, we couple multicomponent (3+1)(3+1)D BFBF theories to formulate an effective field theory for 4d FTOs: $S=\int\frac{A_{I,J}}{2\pi}\,b^{I}\wedge da^{J}$. Stacking $N$ layers of such (3+1)(3+1)D field theories and coupling neighboring 3D “layers” via interlayer $BF$ terms while preserving translational symmetry yields a multicomponent $BF$ theory characterized by a Toeplitz $K$ matrix with integer elements:
@@
Given an iBFBF theory with nearest-neighbor coupling, a natural question arises: under what conditions does nonlocal braiding along the stacking direction emerge? Is there an analog of boundary-mode-induced nonlocal braiding—i.e., Toeplitz braiding— along the stacking direction? To address this, we classify iBFBF theories into two types. The first involves symmetric $K$ matrices, where $C=B^{\mathsf{T}}$ in Eq. (2); in this case, the Toeplitz braiding analysis parallels that of iCS theories [109], except that the particle–particle braiding of Ref. [109] now corresponds to particle–loop braiding in the $BF$ context.
@@
When all “left” ZSMs (LZSMs) localize on one $w$ boundary and “right” ZSMs (RZSMs) on the other, [^4] Here the “left” and the “right” label internal sectors of the SVD rather than spatial sides. either the lower-left or upper-right elements of $K^{-1}$ decay exponentially with system size. Consequently, nontrivial braiding arises when particles and loops occupy specific opposite $w$ boundaries, as illustrated in Figs. 1(a) and 1(b). The corresponding braiding phases are encoded in the lower-left or upper-right elements of $K^{-1}$ [Figs. 1(c) and 1(d)]. In either case, exchanging the particle and loop positions rapidly suppresses the braiding phase upon increasing the size of the $w$ direction. In particular, iBFBF theories with $K$ matrices resembling the nSSH Hamiltonians can exhibit nontrivial Toeplitz braiding in both configurations shown schematically in Figs. 1(a) and 1(b), as the $K$ matrices may host two LZSMs (RZSMs) that localize at distinct $w$ boundaries. The relation between the locations of the LZSMs and RZSMs, the patterns of $K^{-1}$, and the resulting braiding statistics is summarized in Fig. 1(e).
@@
The paper is organized as follows. Section II establishes the braiding structure of multicomponent $BF$ theories and introduces the stacking construction that gives rise to iBFBF field theories, emphasizing why non-diagonal Toeplitz $K$ matrices are essential. Sections III and IV uncover Toeplitz braiding in iBFBF theories whose $K$ matrices inherit the mathematical structures of the HN and nSSH Hamiltonians, respectively. These sections show how SVD sharply exposes boundary zero singular modes and reveal a direct analogy to non-Hermitian directional amplification. Section V closes with a brief summary and outlook.